OpenStudy (anonymous):

Solve the equation. Then check your solution. 4/5+x=3/7 a.13/35 b.-1/2 c.1. 8/12 <--- = one and eight twelfths D. -13/35

OpenStudy (anonymous):

A

OpenStudy (texaschic101):

4/5 + x = 3/7 x = 3/7 -4/5 x = 15/35 - 28/35 x = - 13/35

OpenStudy (texaschic101):

and if you want to check it, sub in - 13/35 in for x in the original equation and see if both sides equal.

OpenStudy (anonymous):

a is wrong

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i submitted my answer and it says a is wrong so it must be d

OpenStudy (anonymous):

probaly

OpenStudy (texaschic101):

it is D

OpenStudy (anonymous):

yep it is thanks

OpenStudy (texaschic101):

yw